OpenStudy (anonymous):

two 10cm diameter metal disks separated by a 0.63mm thick piece of pyrex glass are charged to a potential difference of 1000V. dialectric constant for pyrex is K=4.7. what is the surface charge density on the disks?

OpenStudy (anonymous):

i found that parallel-plate capacitor has \[C=(\epsilon _{0}A)/d\] using that, i get C=(8.85*10^-12*pi*.05^2)/.00063=1.1*10^-10F then Q=CV=1.1*10^-10*1000=1.1*10^-7C then \[\eta=Q/A\] so (1.1*10^-7)/(pi*.05^2)=1.4*10^-5C/m^2

OpenStudy (anonymous):

what i did just doesn't seem right since there should be A LOT of charge per square meter...right?

OpenStudy (ash2326):

I think you are right.